Case 2 The management of a Saudi Manufacturing Company has asked for your assistance in deciding...

Case 2
The management of a Saudi Manufacturing Company has asked for your assistance in deciding whether to continue manufacturing a part or to buy it from an outside supplier. The part, called Tropica, is a component of the company's finished product.
An analysis of the accounting records and the production data revealed the following information for the year ending December 31, 2018.
1. The Machinery Department produced 36,000 units of Tropica.
2. Each Tropica unit requires 10 minutes to produce. Three people in the Machinery Department work full time (2,000 hours per year) to produce Tropica. Each person is paid SR11.00 per hour.
3. The cost of materials per Tropica unit is SR2.00.
4. Manufacturing overhead costs directly applicable to the production of Tropica are: indirect labor, SR5,500; utilities, SR1,300; depreciation, SR1,600; property taxes and insurance, SR1,000. All of the costs will be eliminated if Tropica is purchased.
5. The lowest price for a Tropica from an outside supplier is SR3.90 per unit. Freight charges will be SR0.30 per unit, and a part-time receiving clerk at SR8,500 per year will be required.
6. If Tropica is purchased, the excess space will be used to produce SR10,000 of net income in addition to the savings on the rental of storage space. Currently, the company rents storage space at approximately SR0.60 per unit stored per year. Approximately 6,000 units per year are stored in the rented space.
Required
Should the company make or buy the part (1 mark)?
Answer #1

Total Units = 36,000

Main power rate per hour = 11

Total Hours = 6000

Total Main Power Exp = 11* 6000 = 66000

Cost of material per unit = 2

Total Cost of Material = 2 * 36000 = 72000

Indirect Labor = 5500

Utilities = 1300

Depreciation = 1600

Property Taxes and Insurances = 1000

Rental Storage Space = 6000 * 0.60 = 3600

Total Cost if produced internally = 66000 + 72000 + 5500 + 1300 + 1600 + 1000 + 3600 = 151000

If Tropica produced internally cost is 151000.

If purchased

Purchase price 3.90 per unit

Freight charges 0.30 per unit

Part-time receiving clerk at SR8,500

Income from the use if excess space 10000

Total Purchase Price = 3.90 * 36000 = 140400

Total Freight charges = 0.30 * 36000 = 10800

Total Cost if Purchased = 140400 + 10800 + 8500 - 10000 (Income) = 149700

Total Cost of Making = 151000

Total Cost of Buying = 149700

So company should buy.